OpenStudy (anonymous):

Solve the equation for y: -2x + 4y = 8

OpenStudy (anonymous):

y = 2x + 8 y = 4(2x + 8) y = -2x - 8 none of the above

OpenStudy (anonymous):

-2x +4y =8 -> 4y = 2x + 8 -> y = 1/2x + 2

OpenStudy (anonymous):

none of the above